# Imperative Learning: A Self-supervised Neuro-Symbolic Learning Framework for Robot Autonomy

### v5

@@ 3.2 Formulation
The memory system within the IL framework is a general component that can retain and retrieve information online. Specifically, it can be any structure that is associated with write and read operations to retain and retrieve data (Wang et al., 2021a). A memory can be a neural network, where information is “written” into the parameters and is “read” through a set of math operations or implicit mapping, e.g., a neural radiance fields (NeRF) model (Mildenhall et al., 2021); It can also be a structure with explicit physical meanings such as a map created online, a set of logical rules inducted online, or even a dataset collected online; It can also be the memory system of LLMs in textual form, such as retrieval-augmented generation (RAG) (Lewis et al., 2020), which writes, reads, and manages symbolic knowledge.
@@ 4 Applications and Examples
Additionally, since IL is a self-supervised learning framework consisting of three primary components, we have three kinds of learning types. This includes (A) given known (pre-trained or human-defined) reasoning engines such as logical reasoning, physical principles, and geometrical analytics, robots can learn a logic-, physics-, or geometry-aware neural perception system, respectively, in a self-supervised manner; (B) given neural perception systems such as a vision foundation model (Kirillov et al., 2023), robots can discover the world rules, e.g., traffic rules, and then apply the rules to future events; and (C) given a memory system, (e.g., experience, world rules, or maps), robots can simultaneously update the neural system and reasoning engine so that they can adapt to novel environments with a new set of rules. The five examples will also cover all three learning types.
@@ 5.1.1 Baselines
To evaluate the generalization ability of the proposed iA framework, we select both classic and data-driven methods as the baselines. Specifically, we will compare iA with the popular classic method, weighted A (Pohl, 1970) and the latest data-driven and open-sourced methods including neural A (Yonetani et al., 2021b), focal search with path probability map (FS + PPM) and weighted A with correction factor (WA + CF) (Kirilenko et al., 2023).
@@ 5.1.5 Qualitative Performance
Figure 10: Real-world experiment for local path planning using iPlanner with a legged robot. The red curve indicates the robot’s trajectory from right to left, beginning inside a building and then navigating to the outdoors. The robot follows a series of waypoints (blue) and plans in different scenarios marked by green boxes including (A) passing through doorways, (B, D, E) circumventing both static and dynamic obstacles, and (B, F) ascending and descending stairs.
@@ 5.3.1 Baselines
To demonstrate the effectiveness of the IL framework for jointly optimizing the data-driven IMU model and MPC, we select four baselines including (1) IMU+MPC: classic IMU integrator with a regular Diff-MPC; (2) IMU++MPC: data-driven IMU model with a regular Diff-MPC; (3) IMU+MPC+: classic IMU integrator with a Diff-MPC with learnable moment of inertia (MOI); and (4) iMPC (ours): IMU+ with MPC+ trained with IL framework. Specifically, “IMU+” refers to the data-driven IMU denoising and integration model from (Qiu et al., 2024).
@@
5.3.2 Implementation Details
Due to efficiency considerations, we use a lightweight network for the data-driven IMU model, which includes a 2-layer multi-layer perceptron (MLP) as the encoder to get the feature embeddings from both the accelerometer and the gyroscope. Two decoders are implemented each comprising two linear layers to address the specific corrections needed for the accelerometer and the gyroscope. These decoders are dedicated to accurately predicting the necessary adjustments for each sensor type, thereby enhancing the overall precision of the IMU readings in dynamic conditions. For both the encoder and the decoder, we use GELU as the activation function for its effectiveness in facilitating smoother non-linear transformations.
@@ 5.3.4 Metrics
5.3.5 Robustness to Initial Conditions
Figure 13: Control performance of iMPC under different levels of wind disturbance.
@@ 5.3.5 Robustness to Initial Conditions
Table 10: UAV attitude control performance under impulse wind disturbance for different speeds.
@@
5.3.6 Robustness to Wind Disturbance
We next evaluate the control performance under the wind disturbance to validate the robustness of the proposed approach. Specifically, we inject the wind disturbance by applying an external force $F=\frac{1}{2}\cdot C_{d}\cdot\rho\cdot A\cdot V^{2}$ and torque $\tau=r\times F$ to the nominal system based on the drag equation, where $C_{d}$ is the drag coefficient, $\rho$ is the air density, $A$ is the frontal area of the object facing the wind, $V$ is the wind speed, $r$ is the length of lever arm. Hence, for a fixed UAV configuration, different wind speeds $V$ lead to different disturbance levels. We conduct two sets of experiments: in the first set, we add impulse wind gust to the UAV in the direction of opposite heading during the hover with varying wind speed $10/15/20/45\ \mathrm{m}/\mathrm{s}^{2}$; in the second set, a step wind is added to the system in the same direction during hover and lasts for 0.3 s  for different wind speed $10/15/20/35\ \mathrm{m}/\mathrm{s}^{2}$. The external wind starts at 0.2 s  after entering a steady hover.
@@ 5.3.6 Robustness to Wind Disturbance
Figure 14: iSLAM tracks the robot’s trajectory (left) in real-time and simultaneously performs a dense reconstruction (right 1-4).
Table 11: UAV attitude control performance under step wind disturbance for different speeds.
@@
Detailed comparisons with other approaches are presented in Table 10 and Table 11 for the two scenarios with different disturbances, respectively. It is evident that iMPC leads to better attitude control performance with smaller SSE and a faster ST. Particularly compared to IMU+MPC and IMU++MPC, iMPC reduces noises with a more accurate attitude state prediction. Additionally, iMPC further reduces the IMU prediction error in terms of RMSE. Therefore, even in the presence of significant external wind disturbances, iMPC can still control the vehicle attitude robustly. This further demonstrates that the IL framework can simultaneously enhance both the MPC parameter learning and IMU denoising and prediction.
@@ 5.4.1 Datasets
To ensure thorough evaluation, we adopt three widely-used datasets, including KITTI (Geiger et al., 2013), EuRoC (Burri et al., 2016), and TartanAir (Wang et al., 2020). They have diverse environments and motion patterns: KITTI incorporates high-speed long-range movements in driving scenarios, EuRoC features aggressive motions in indoor environments, and TartanAir offers challenging environments with various illuminations and moving objects.
5.4.2 Metrics
Following prior works (Qiu et al., 2022), we use the widely-used absolute trajectory error (ATE), relative motion error (RME), and root mean square error (RMSE) of rotational and transnational drifts as evaluation metrics.
5.4.3 Accuracy Evaluation
We first assess the localization accuracy of iSLAM on the KITTI dataset, which has been widely used in previous works on various sensor setups. To facilitate a fair comparison, in Table 12, we evaluate our standalone VO component against existing VO networks, including DeepVO (Wang et al., 2017b), UnDeepVO (Li et al., 2018), TartanVO (Wang et al., 2021b), and DROID-SLAM (Teed and Deng, 2021), and compare the full iSLAM to other learning-based visual-inertial methods, which are Wei et al. (2021), Yang et al. (2022a), and DeepVIO (Han et al., 2019). Sequences 00 and 03 are omitted in our experiment since they lack completed IMU data. Notably, some methods such as DeepVO and (Yang et al., 2022a) were supervisedly trained on KITTI, while our iSLAM was self-supervised. Even though, iSLAM outperforms all the competitors. Additionally, it is noteworthy that our base model, TartanVO, doesn’t exhibit the highest performance due to its lightweight design. Nevertheless, through imperative learning, we achieve much lower errors with a similar model architecture. Figure 14 shows the trajectory and reconstruction results on sequence 05.
@@ 5.4.3 Accuracy Evaluation
Table 14: Relative Motion Error (RME) on two environments of TartanAir dataset. ORB-SLAM2 (Mur-Artal and Tardós, 2017), TartanVO (Wang et al., 2021b), and AirVO (Xu et al., 2023) are visual methods; OKVIS (Leutenegger et al., 2015), ORB-SLAM3 (Campos et al., 2021), and our iSLAM are visual-inertial methods. “✗” represents the method lost tracking or drifted too far away.
OKVIS ORB-SLAM2 ORB-SLAM3 TartanVO AirVO iSLAM (ours)
$r_{\text{rel}}$ $t_{\text{rel}}$ $r_{\text{rel}}$ $t_{\text{rel}}$ $r_{\text{rel}}$ $t_{\text{rel}}$ $r_{\text{rel}}$ $t_{\text{rel}}$ $r_{\text{rel}}$ $t_{\text{rel}}$ $r_{\text{rel}}$ $t_{\text{rel}}$
Soulcity Hard P000 1.523 9.628 0.068 1.097 0.185 4.469 0.412 10.134 0.064 3.945
P001 1.257 10.837 0.204 3.442 0.202 4.147 0.071 3.793
P002 1.713 7.596 0.147 1.663 0.069 0.575 0.176 2.081 0.334 4.370 0.066 1.898
P003 1.119 10.520 0.263 5.321 0.258 4.945 0.090 4.776
P004 1.562 11.682 0.167 2.361 0.174 3.391 0.262 4.925 0.082 4.569
P005 0.096 1.333 0.208 4.482 0.526 15.363 0.079 4.227
P008 1.316 9.621 0.112 2.782 0.101 2.862 0.146 3.602 0.050 3.258
P009 1.557 10.283 0.134 2.605 0.194 4.376 0.188 4.427 0.057 3.885
Ocean Hard P000 1.962 12.777 0.129 1.993 0.045 1.723
P001 0.098 1.622 0.157 3.750 0.633 13.294 0.059 3.191
P002 1.889 16.820 0.180 4.004 0.671 14.481 0.060 3.621
P003 2.649 19.019 0.166 4.528 1.140 30.188 0.056 3.978
P004 0.332 3.231 0.152 1.658 0.088 0.644 0.131 1.525 0.354 5.500 0.055 1.280
P005 1.181 4.395 0.212 5.236 0.090 1.636 0.322 4.108 0.041 1.399
P006 0.819 12.810 0.149 3.305 0.898 17.882 0.051 2.859
P007 0.163 3.842 1.169 23.189 0.058 3.575
P008 1.594 13.010 0.102 1.228 0.146 2.862 0.700 12.625 0.052 2.292
P009 1.809 23.033 0.159 6.238 1.224 28.326 0.055 5.572
Avg 1.507 12.205 - - - - 0.172 3.709 0.699 14.955 0.061 3.325
5.4.4 Robustness Assessment
In this section, we evaluate the robustness of iSLAM against other competitors, including the widely-used ORB-SLAM2 (Mur-Artal and Tardós, 2017), ORB-SLAM3 (Campos et al., 2021), and a new hybrid method AirVO (Xu et al., 2023). Two “Hard” level testing environments in TartanAir (Wang et al., 2020) are used, namely Ocean and Soulcity. The Ocean environment has dynamic objects such as fishes and bubbles, while the Soulcity features complex lighting with rainfall and flare effects. Their challenging nature results in many failures of other methods. As shown in Table 14, on 18 testing sequences, ORB-SLAM2 failed on 11, ORB-SLAM3 failed on 7, and AirVO failed on 6. In contrast, iSLAM accurately tracked all the sequences, yielding the best robustness.
Figure 15: (a) The ATE of our VO and PVGO w.r.t. number of imperative iterations. (b) The decrease in error percentage. The error before imperative learning is treated as 100%. The ATE is used for VO and PVGO to calculate the percentage, while the relative displacement error is used for IMU. The solid lines are the mean values on all sequences we used in KITTI, EuRoC, and TartanAir, while the transparent regions are the variances.

### v6

@@ 3.2 Formulation
The memory system within the IL framework is a general component that can retain and retrieve information online. Specifically, it can be any structure that is associated with write and read operations to retain and retrieve data (Wang et al., 2021a). A memory can be a neural network, where information is “written” into the parameters and is “read” through a set of math operations or implicit mapping, e.g., a neural radiance fields (NeRF) model (Mildenhall et al., 2021); It can also be a structure with explicit physical meanings such as a map created online, a set of logical rules inducted online, or even a dataset collected online; It can also be the memory system of LLMs in textual form, such as retrieval-augmented generation (RAG) (Lewis et al., 2020), which writes, reads, and manages symbolic knowledge. Hyperdimensional computing (Kleyko et al., 2022, 2023) also presents a promising direction for memory modeling, though it falls outside the scope of this work. In this article, we will explore several customized memory structures tailored to specific applications, as detailed in Section 4.
@@ 3.3.2 Implicit Differentiation
Discussion on theoretical optimality and sufficient conditions.
The guarantee of feasible solutions depends on the assumptions and conditions of the objective functions. For example, in unrolled differentiation (Section 3.3.1) and implicit differentiation (Section 3.3.2), if the LL function $L$ in (1b) is twice differentiable and strongly convex with respect to $\phi$ with a unique solution, we expect to achieve first-order stationarity and global optimality for non-convex and convex UL objective functions, respectively.
The optimality for generic non-convex LL problems remain an open challenge and are sometimes NP-hard (Dempe et al., 2015) in BLO theory. This difficulty arises because non-uniqueness in the LL problem prevents the implicit differentiation theorem from guaranteeing the existence of implicit differentiability of $\phi^{*}$ with respect to the UL variables $\psi$. A weaker condition for establishing optimality is the Polyak-Lojasiewicz (PL) condition (Shen and Chen, 2023), which allows for an overall non-convex landscape while maintaining certain convexity-like structures. Such structures have been observed in overparameterized neural networks and can guarantee an $\epsilon$-approximate solution to the BLO problem (Liu et al., 2020a). We will show more detailed theoretical guarantees and their assumptions in the next section with respect to the types of LL optimization. To ensure the language remaining accessible and not overly theoretical, we will only provide summarized results and provide references to their sources. In this way, we hope it will be helpful for readers to easily find more relevant formal results when needed.
@@ 4 Applications and Examples
Additionally, since IL is a self-supervised learning framework consisting of three primary components, we have three kinds of learning types. This includes (A) given known (pre-trained or human-defined) reasoning engines such as logical reasoning, physical principles, and geometrical analytics, robots can learn a logic-, physics-, or geometry-aware neural perception system, respectively, in a self-supervised manner; (B) given neural perception systems such as a vision foundation model (Kirillov et al., 2023), robots can discover the world rules, e.g., traffic rules, and then apply the rules to future events; and (C) given a memory system, (e.g., experience, world rules, or maps), robots can simultaneously update the neural system and reasoning engine so that they can adapt to novel environments with a new set of rules. Note that if one of the modules has a good initialization, all the three modules can be further finetuned simultaneously. The five examples will cover all three learning types.
@@ 4.2 First-order Optimization
Discussion on theoretical guarantee and optimality
To achieve theoretical guarantees, one sufficient condition is that the LL function $L$ in (1b) is twice differentiable and strongly convex respect to $\phi$. The UL function $U$ can be possibly non-convex. Based on (Ji et al., 2021, Theorem 1), the proposed first-order method, which solves the linear system in (31) and uses $\hat{\bm{\mu}}$ and $\hat{\bm{\nu}}$ of the solutions ${\bm{\mu}}^{*}$ and ${\bm{\nu}}^{*}$, can achieve the first-order $\epsilon$-accurate stationary optimality, i.e., $\|\nabla_{\psi}\|\leq\epsilon$, if we use $\log\frac{1}{\epsilon}$-level steps for the approximation in (32). Moreover, when the UL function $U$ in (19) is strongly convex with respect to $\theta$ and $\gamma$, we can further show that the obtained solution achieves the global optimality, based on (Ji and Liang, 2021, Theorem 9).
@@ 4.3 Constrained Optimization
Discussion on theoretical guarantee and optimality
When the Hessian matrices $\frac{\partial^{2}L}{\partial^{2}\phi}$ and $\frac{\partial^{2}\xi}{\partial^{2}\phi}$ of the LL function $L$ and the constraint function $\xi$ are invertible, we can guarantee the matrix $H$ in (37) is invertible. One sufficient condition for such invertibility is that $L$ and $\xi$ are strongly convex. Motivated by the approximate Karush-Kuhn-Tucker (KKT) condition (Andreani et al., 2010), which is characterized as an optimality condition for nonlinear program, we can use the residual function in (Yao et al., 2024b, Eq. (15)) as an optimality measurement. As a result, the optimality can be established and guaranteed under these conditions.
@@ 4.4 Second-order Optimization
Discussion on theoretical guarantee and optimality
The theoretical guarantee is similar to that for the first-order methods in Section 4.2, as long as the second-order LL optimizers such as Gauss-Newton and LM approximate $\mu^{*}$ and $\nu^{*}$ to the target accuracy $\epsilon$. For example, if the LL problem is strongly convex, the second-order optimization can find an $\epsilon$-accurate solution with an accelerated linear convergence rate (Ji and Liang, 2021, Theorem 9 and 10).
@@ 5.1.1 Baselines
To evaluate the generalization ability of the proposed iA framework, we select both classic and data-driven methods as the baselines. Specifically, we will compare iA with the popular classic method, weighted A (Pohl, 1970) and the latest imitation learning-based methods including neural A (Yonetani et al., 2021b), focal search with path probability map (FS + PPM) and weighted A with correction factor (WA + CF) (Kirilenko et al., 2023).
@@ 5.1.9 Quantitative Performance
5.1.10 Efficiency Analysis
We evaluate iPlanner’s efficiency in Matterport3D. The robot follows 20 manual waypoints for global guidance but autonomously plans feasible paths and avoids local obstacles. As shown in Figure LABEL:fig:computing-time, iPlanner achieves an average 3×3$\times$3 × speedup over MPP. The executed trajectory is shown in Figure LABEL:fig:trajectory.
@@ 5.1.11 Real-world Tests
Figure 10: An example of real-time planning using iPlanner. (a) shows the comparison of planning time. (b) shows the trajectory executed by iPlanner and the outline of the environment.
@@ 5.2.1 Datasets
Figure 11: Real-world experiment for local path planning using iPlanner with a legged robot. The red curve indicates the robot’s trajectory, beginning inside a building and then navigating to the outdoors. The robot follows a series of waypoints (blue) and plans in different scenarios marked by green boxes including (A) passing through doorways and avoid dynamic obstacles, (B) ascending stairs, and (C) circumvent outdoor static obstacles.
@@ 5.2.4 Evaluation Results
5.2.5 Efficiency Analysis
Due to the complicated BLO, the training time for iLogic is about 40% longer than the supervised learning “ResNet + NLM”, and this extra cost is well justified by the significant performance gain in Table 6 and Table 7. However, there is no additional computational cost in inference, with the system achieving a speed of approximately 19.78 frames/second on an RTX 3090 GPU.
@@ 5.3.1 Baselines
To demonstrate the effectiveness of the IL framework for jointly optimizing the data-driven IMU model and MPC, we select four baselines including (1) IMU+MPC: classic IMU integrator with a regular Diff-MPC; (2) IMU++MPC: data-driven IMU model with a regular Diff-MPC; (3) IMU+MPC+: classic IMU integrator with a Diff-MPC with learnable moment of inertia (MOI); and (4) iMPC (ours): IMU+ with MPC+ trained with IL framework. Specifically, “IMU+” refers to the data-driven IMU denoising and integration model from (Qiu et al., 2024). Additionally, for the wind disturbance test, we select an RL approach with a commonly used approach, proximal policy optimization (PPO) (Schulman et al., 2017) as the baseline.
5.3.2 Implementation Details
Due to efficiency considerations, we use a lightweight network for the data-driven IMU model, which includes a 2-layer multi-layer perceptron (MLP) as the encoder to get the feature embeddings from both the accelerometer and the gyroscope. Two decoders are implemented each comprising two linear layers to address the specific corrections needed for the accelerometer and the gyroscope. These decoders are dedicated to accurately predicting the necessary adjustments for each sensor type, thereby enhancing the overall precision of the IMU readings in dynamic conditions. For both the encoder and the decoder, we use GELU as the activation function for its effectiveness in facilitating smoother non-linear transformations. For RL, we take pre-integrated IMU measurements as the observation, output UAV action (thrust and torques) in OpenAI gym (Brockman et al., 2016), and designed the reward to minimize the difference between the current and target UAV pose.
@@ 5.3.5 Robustness to Initial Conditions
Figure 14: Control performance of iMPC and PPO under different levels of wind disturbance.
@@
5.3.6 Robustness to Wind Disturbance
We next evaluate the control performance under the wind disturbance to validate the robustness of the proposed approach. Specifically, we inject the wind disturbance by applying an external force $F=\frac{1}{2}\cdot C_{d}\cdot\rho\cdot A\cdot V^{2}$ and torque $\tau=r\times F$ to the nominal system based on the drag equation, where $C_{d}$ is the drag coefficient, $\rho$ is the air density, $A$ is the frontal area of the object facing the wind, $V$ is the wind speed, $r$ is the length of lever arm. Hence, for a fixed UAV configuration, different wind speeds $V$ lead to different disturbance levels. We conduct two sets of experiments: in the first set, we add impulse wind gust to the UAV in the direction of opposite heading during the hover with varying wind speed $10/15/20/45\ \mathrm{m}/\mathrm{s}$; in the second set, a step wind is added to the system in the same direction during hover and lasts for 0.3 s  for different wind speed $10/15/20/35\ \mathrm{m}/\mathrm{s}$. The external wind starts at 0.2 s  after entering a steady hover. During RL (PPO) training, we apply external wind disturbances of up to $10\mathrm{m}/\mathrm{s}$. In contrast, our hybrid approach with MPC provides robustness without requiring exposure to a wide range of wind conditions during training.
Table 10: UAV attitude control performance under impulse wind, where ‘✗’ means losing control at that wind speed.
@@ 5.3.6 Robustness to Wind Disturbance
Table 11: UAV attitude control performance under step wind, where ‘✗’ means losing control at that wind speed.
@@
Detailed comparisons with other approaches are presented in Table 10 and Table 11 for the two scenarios with different disturbances, respectively. It is evident that iMPC leads to better attitude control performance with smaller SSE and a faster ST. Particularly compared to IMU+MPC and IMU++MPC, iMPC reduces noises with a more accurate attitude state prediction. Additionally, iMPC further reduces the IMU prediction error in terms of RMSE. Therefore, even in the presence of significant external wind disturbances, iMPC can still control the vehicle attitude robustly. This demonstrates that the IL framework can simultaneously enhance both the MPC parameter learning and IMU denoising and prediction. Compared to the RL baseline PPO, iMPC demonstrates significantly better performance, especially under wind disturbances that were not encountered during training. While PPO performs similarly under wind conditions seen during training, as shown in Figure LABEL:fig:wind-impulse and Figure LABEL:fig:wind-step, it completely loses control of the UAV when faced with out-of-distribution disturbances. These results further highlight the robustness of the IL framework, benefiting from a hybrid approach that integrates prior physical knowledge.
5.3.7 Efficiency Analysis
Compared to a conventional approach of IMU+MPC, iMPC achieves similar efficiency because the MPC dominates the total computation time and the overhead from the learning-based IMU inference is negligible. In fact, iMPC runs at 50 Hz control frequency with a 200 Hz IMU measurement, which is at the same level as standard MPC. On the other hand, when compared to RL, although the inference time of iMPC is longer as expected due to the additional requirement of solving an online optimization, such physics-based optimization also brings prior knowledge into the training process, thereby overcoming the typical low sampling efficiency issue of RL. In fact, the experiment shows the PPO needs an extra 221.28% time for training due to the low sampling efficiency.
@@ 5.4.1 Datasets
To ensure thorough evaluation, we adopt two widely-used datasets, including KITTI (Geiger et al., 2013) and EuRoC (Burri et al., 2016). They have diverse environments and motion patterns: KITTI incorporates high-speed long-range movements in driving scenarios, while EuRoC features aggressive motions in indoor environments.
@@
5.4.2 Metrics
Following prior works (Qiu et al., 2022), we use the widely-used absolute trajectory error (ATE), relative motion error (RME), and root mean square error (RMSE) of rotational and translational drifts as evaluation metrics.
5.4.3 Accuracy Evaluation
We first assess the localization accuracy of iSLAM on the KITTI dataset, which has been widely used in previous works on various sensor setups. To facilitate a fair comparison, in Table 12, we evaluate our standalone VO component against existing VO networks, including DeepVO (Wang et al., 2017b), UnDeepVO (Li et al., 2018), TartanVO (Wang et al., 2021b), and DROID-SLAM (Teed and Deng, 2021), and compare the full iSLAM to other learning-based visual-inertial methods, which are Wei et al. (2021), Yang et al. (2022a), and DeepVIO (Han et al., 2019). Sequences 00 and 03 are omitted in our experiment since they lack completed IMU data. Notably, some methods such as DeepVO and (Yang et al., 2022a) were supervisedly trained on KITTI, while our iSLAM was self-supervised. Even though, iSLAM outperforms all the competitors. Additionally, it is noteworthy that our base model, TartanVO, doesn’t exhibit the highest performance due to its lightweight design. Nevertheless, through imperative learning, we achieve much lower errors with a similar model architecture.
@@ 5.4.3 Accuracy Evaluation
Figure 15: (a) The ATE of our VO and PVGO w.r.t. number of imperative iterations. (b) The decrease in error percentage. The error before imperative learning is treated as 100%. The ATE is used for VO and PVGO to calculate the percentage, while the relative displacement error is used for IMU. The solid lines are mean values while the transparent regions are the variances.